# Generative AI in Transportation Planning: A Survey

### v4

@@ Abstract
Figure 1: Generative AI is revolutionizing transportation planning through advanced language analysis and interdisciplinary integration capabilities.
@@ 1 Introduction
Structure. Section 2 provides background knowledge, followed by our taxonomy in Section 3. Sections 4 and 5 present transportation and computational perspectives, respectively. Section 6 discusses challenges and future directions, and Section 7 concludes the survey.
2 Preliminaries
@@ 2.1 Transportation Planning
2.2 Evolution of AI in Transportation
The latest evolution—generative AI—represents a paradigm shift toward highly autonomous systems. Generative models leverage large-scale, fine-grained datasets, including real-time sensor inputs and external sources like social media or weather forecasts [69, 70]. These models autonomously generate solutions, simulate traffic scenarios, and optimize infrastructure designs without requiring explicit programming for each task [8, 71]. However, the integration of existing expert knowledge remains crucial in guiding these AI systems. Domain expertise is essential for defining objectives, validating outputs, and ensuring alignment with established transportation modeling principles. Expert insights help refine AI-driven processes by providing a clear direction, especially in complex tasks like multimodal demand forecasting or infrastructure planning, where understanding system nuances is critical. Generative AI systems, when combined with expert knowledge, enable rapid execution and real-time decision-making, accelerating tasks such as traffic rerouting during emergencies, infrastructure design under various constraints, and automated reporting for public engagement.
2.3 Generative AI Models for Transportation
@@
Generative AI’s impact is particularly notable in the following advancements:
  • Scenario Simulation and Evaluation: Generative AI enables the creation of synthetic transportation scenarios to evaluate policy impacts, infrastructure designs, and environmental consequences under diverse conditions [87]. These capabilities support data-driven decision-making for long-term urban planning and emergency response preparation.
  • Traffic Prediction and Decision-Making: Advanced AI models predict congestion patterns, vehicle flows, and multimodal interactions with high accuracy. By integrating real-time data with historical trends, generative models optimize traffic signals, routing strategies, and operational efficiency [88, 89].
  • Data Fusion and Enhancement: Generative AI improves data integration by synthesizing missing or sparse data, enabling better analytics and visualization. This is critical for applications such as safety analysis, where detailed accident simulations require high-resolution data [49, 50].
  • Public Engagement and Sentiment Analysis: Generative AI processes public feedback from surveys, social media, and other sources to assess community needs and preferences. These insights guide policy development and foster inclusive decision-making [90].
Recent studies have also emphasized the role of LLMs in supporting generative AI tasks. LLMs enhance scenario generation, explainability, and interactive planning through advanced text generation and reasoning capabilities [91]. For instance, LLMs can generate detailed descriptions of infrastructure projects, analyze public sentiment on transportation policies, and simulate stakeholder discussions, bridging the gap between technical outputs and human decision-making processes.
3 Taxonomy on Generative AI for Transportation Planning
To systematically classify the role of generative AI in transportation planning, we introduce a taxonomy organized into two major categories: (1) Core Transportation Applications and (2) Generative AI Methodologies. This taxonomy aligns computational innovations with practical transportation goals, such as improving mobility efficiency, promoting sustainability, and fostering equitable infrastructure development.
3.1 Core Transportation Applications
@@
Demand Forecasting applies generative AI to predict travel demand patterns across multimodal systems. Models can fine-tune origin-destination (O-D) matrices to align with observed traffic counts [109, 7], estimate the adoption of shared mobility services [110, 111, 112], and simulate long-term demand shifts due to demographic or economic changes [70, 52, 113, 88, 114]. These forecasts provide insights into system bottlenecks, enabling planners to balance travel loads across transportation networks effectively.
@@
3.2 Generative AI Methodologies
Benchmark Datasets form the foundation for evaluating generative AI performance in tasks like O-D calibration, traffic simulation, and infrastructure planning. Examples include LargeST [124] for traffic analysis and traffic assignment datasets [125]. These datasets provide high-resolution inputs and ground-truth benchmarks to validate AI-driven insights. However, building comprehensive knowledge bases for these models remains a significant challenge. Particularly for forecasting tasks, where uncertainty and dynamic variables play key roles, such as: population growth, land-use changes, and evolving travel behaviors. Generative AI requires diverse, harmonized data inputs, such as historical traffic data, multimodal interactions, and prior model outcomes, but integrating these into a machine-readable format is time-intensive and complex. The lack of standardization, interoperability, and real-time updates further complicates the process, creating a bottleneck in AI adoption. Future advancements must focus on automated dataset synthesis, standardized modeling frameworks, and adaptive real-time updates to ensure that generative AI can address the intricate and evolving demands of transportation systems effectively.
Data Preprocessing Strategies ensure clean, balanced, and representative inputs for AI models. Transportation datasets often require manual or automated annotation of traffic patterns [126], infrastructure attributes [127], or user feedback [128]. Generative AI methods, such as LLM-based automated labeling [129, 130], address annotation challenges while preserving data quality. Data augmentation techniques, such as synthetic O-D matrix generation [131, 132], further expand datasets to simulate rare or extreme transportation scenarios.
Fine-Tuning Techniques enable generative AI to adapt to transportation-specific tasks efficiently. Methods like Low-Rank Adaptation (LoRA) [133] optimize model performance while minimizing computational demands. For complex tasks like O-D calibration, Chain-of-Thought prompting [115] enhances the model’s reasoning ability to align O-D matrices with observed traffic counts.
@@ 3.2 Generative AI Methodologies
3.3 Summary of the Taxonomy
This taxonomy systematically organizes generative AI applications and methodologies for transportation planning. By connecting AI advancements with core planning objectives—such as scenario generation, demand forecasting, and sustainability optimization—it provides a comprehensive framework to guide research and practical implementation. The methodologies outlined ensure that generative AI models are reliable, scalable, and capable of addressing the unique challenges inherent to modern transportation systems.
@@ 4.3 Generative Tasks for Data Synthesis and Scenario Generation
4.4 Simulation Tasks for Traffic Dynamics and Mixed-Autonomy Systems
@@
Generative AI-driven simulations are particularly valuable for studying mixed-autonomy systems, where human-driven vehicles (HVs) and autonomous vehicles (AVs) coexist. For example, frameworks like CoMAL [115] use LLM agents to simulate collaborative strategies for traffic merging, lane switching, or intersection coordination. These simulations optimize system-level performance by dynamically assigning roles (e.g., leader, follower) to AVs, reducing congestion and stop-and-go waves.
@@
  • Computational cost: Another critical challenge is the computational cost associated with running high-resolution, real-time simulations at scale. Generative AI models integrated into digital twins or predictive traffic systems require significant computational resources to process diverse data streams, such as real-time sensor feeds, GPS traces, and dynamic O-D matrices [199]. While digital twins offer an opportunity to mirror real-world transportation systems for testing interventions, their reliance on generative AI magnifies the resource demands, particularly for urban-scale or multimodal networks. Developing lightweight and computationally efficient models is crucial to for scalability constraints.
@@ 5.1 Benchmark Datasets
5.2 Dataset Preparation Strategies
@@
5.3 Broad Source of Dataset Collection
@@
Ensuring Data Diversity and Representativeness. To ensure generalizability, transportation datasets must represent diverse geographies, demographic groups, and environmental conditions. For instance, combining urban, suburban, and rural datasets enables generative AI models to address transportation challenges across different spatial contexts. In addition, datasets incorporating varied socioeconomic conditions, such as underserved communities or emerging economies, help mitigate biases and promote equitable solutions. The NYC Taxi Dataset [249], for example, combines demographic and geographic attributes to simulate ride demand across neighborhoods with diverse income levels. Similarly, integrating global data sources, such as OpenStreetMap and GTFS feeds[264], ensures that generative AI models remain adaptable to multimodal networks across regions.
Data Augmentation for Generative Tasks. Data augmentation techniques enhance dataset size and diversity, especially when real-world data is scarce or incomplete. For traffic forecasting, techniques such as synthetic trajectory generation and perturbation modeling can create realistic yet varied travel patterns. In infrastructure optimization, generative models can simulate alternative urban layouts, enabling planners to evaluate multiple design scenarios under different constraints, such as cost, population density, and environmental impact. For sentiment analysis, data augmentation methods such as paraphrasing and noise injection can expand labeled datasets, ensuring that generative AI models remain robust to linguistic variations and ambiguities. This is particularly relevant for analyzing multilingual user feedback or regional dialects in transportation systems.
@@
5.4 Fine-Tuning Generative AI for Transportation Planning.
This subsection explores the fine-tuning of generative AI models for transportation planning applications, using Origin-Destination (O-D) Demand Calibration as an example. O-D calibration refers to the process of estimating travel demand between different geographic locations by adjusting O-D matrices to match observed traffic counts. This task is fundamental in transportation systems, as accurate demand estimation directly influences infrastructure planning, traffic control strategies, and network optimization. O-D matrices are critical inputs for modeling transportation demand and evaluating policy interventions, as they help planners understand mobility patterns and design efficient transportation systems. Specifically, we employ the Open Traffic Intelligence (OpenTI) dataset [7], which provides detailed traffic data for O-D calibration tasks. O-D calibration involves adjusting O-D matrices to align with observed traffic counts, a critical problem in transportation systems where accurate demand estimation directly impacts infrastructure planning, traffic control, and network optimization.
Domain-Specific Dataset The OpenTI dataset [7] is an open-source resource specifically designed to support traffic analysis and calibration tasks. It includes traffic flow observations, O-D matrices, road network structures, and simulated travel demands. OpenTI’s detailed data structure provides both: (1) Input: Raw OD matrices and observed traffic counts across multiple time intervals. (2) Output: A series of calibrated OD matrices that better align with real-world traffic conditions. OpenTI captures diverse scenarios, including traffic demand variations during peak and off-peak hours, enabling fine-tuned generative AI models to improve the accuracy and adaptability of OD demand estimation.
Fine-Tuning Process for LLMs. Fine-tuning involves adapting a pre-trained LLM (e.g., Llama3 [266]) to the specific requirements of legislative summarization. The procedure includes the following steps:
@@
    Data Preprocessing: Preprocessing OpenTI data ensures consistency between inputs (raw OD matrices and observed traffic counts) and outputs (calibrated OD matrices). The steps include: (1) Normalization: Scaling OD demand values to ensure uniform ranges across datasets. (2) Matrix Formatting: Aligning OD matrices with observed traffic counts to maintain input-output coherence. (3) Noise Reduction: Filtering out inconsistencies or redundant data points to streamline the training process. The preprocessed data pairs provide well-structured examples for training generative AI models to produce calibrated OD matrices.
@@
    Fine-Tuning Setup: To fine-tune the generative AI model efficiently, parameter-efficient techniques such as Low-Rank Adaptation (LoRA) [133] or prefix-tuning [267] are applied. These techniques modify only a minimal subset of model parameters, reducing computational overhead while maintaining high accuracy. Hyperparameters, including learning rate, batch size, and iteration count, are optimized using evaluation metrics relevant to OD calibration: (1) Relative Error (RE): Measures the difference between the predicted and observed traffic counts. (2) Root Mean Square Error (RMSE): Evaluates the overall accuracy of the calibrated OD matrices.
@@
    Training Process: Training is conducted on GPU clusters for performance optimization, using methods like gradient accumulation [268] and mixed-precision training [269] to efficiently manage large-scale OD matrices. The generative AI model iteratively minimizes the loss function by comparing calibrated OD matrices with ground-truth matrices provided in the OpenTI dataset. Please note that the OpenTI dataset provides a series of ground-truth matrices, which allows for iteratively training for one single scenario. Regular validation is performed on unseen subsets of the data to prevent overfitting and ensure robust generalization across different traffic scenarios, such as varying road conditions, demand peaks, and network configurations.
@@
    Prompt Engineering for OD Calibration: Prompt engineering guides the generative AI model in focusing on critical OD calibration tasks, such as adjusting demand to match observed traffic counts while preserving network flow constraints. Prompts are structured to clarify the task and specify output expectations.
@@
    These prompts ensure that the model outputs calibrated OD matrices that are both accurate and practically usable in traffic optimization and planning tasks.
Expected Generative AI Outputs After fine-tuning on the OpenTI dataset, the generative AI model is expected to perform effectively across several OD calibration tasks:
@@
5.5 Inference with LLMs: Zero-Shot In-Context Learning
This subsection demonstrates the application of Zero-Shot Learning (ZSL) using Large Language Models (LLMs) to estimate city-level transportation accessibility. Inspired by the study "What Can LLM Tell Us About Cities?" [134], we show how LLMs can infer insights about cities without requiring task-specific fine-tuning or labeled datasets. Specifically, we focus on predicting public transportation accessibility metrics, such as the People Near Transit (PNT) score.
Zero-Shot Learning (ZSL) in Urban Studies. ZSL leverages the pre-trained capabilities of LLMs to solve downstream tasks with only textual prompts and no additional training [270]. In city studies, such as transportation accessibility, ZSL enables the extraction of meaningful insights from LLMs by asking well-crafted prompts, even when ground-truth data is unavailable. This makes ZSL an efficient and scalable solution for analyzing cities that lack comprehensive datasets.
Prompt Engineering for ZSL. In the case of estimating transportation accessibility, the goal is to guide the LLM to predict metrics like PNT without explicit training. Effective prompts incorporate contextual details about the city, transportation infrastructure, and government policies. Below is an example prompt:

### v5

@@ Abstract
Figure 1: Generative AI is revolutionizing transportation planning through advanced language analysis and interdisciplinary integration capabilities. (Content Warning: This paper may contain offensive content generated by LLMs)
@@ 1 Introduction
Structure. As shown in Figure 2, Section 2 provides background knowledge. Section 3 present classical transportation planning tasks and computational perspectives from GenAI, respectively. Section 4 shows some application-specific applications to dataset and experimental results. Section 5 discusses challenges and future directions, and Section 6 concludes the survey.
{forest} for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=16em,font= ,, where level=2text width=22em,font= ,, where level=3text width=13em,font= ,, where level=4text width=7em,font= ,, [ The Generative AI in
Transportation Planning, [ Introduction (Sec. 1), ], [ Background (Sec. 2) [ Transportation Planning (Sec. 2.1) ], [ GenAI Methodologies (Sec. 2.2) ], [ Evolution of AI in Transportation (Sec. 2.3) ] ], [ Classical Planning Functions and
Modern Transformations (Sec. 3) [ Descriptive Tasks (Sec. 3.1) ], [ Predictive Tasks (Sec. 3.2) ], [ Generative Tasks (Sec. 3.3) ], [ Simulation Tasks (Sec. 3.4) ], [ Trustworthiness (Sec. 3.5) ] ], [ Technical Foundations for GenAI
in Transportation Planning (Sec. 4) [ Dataset Preparation (Sec. 4.1) ], [ Fine-tuning GenAI Models (Sec. 4.2) ], [ Zero-shot Inference with GenAI (Sec. 4.3) ], [ Few-Shot Inference with GenAI (Sec. 4.4) ], [ Other Techniques Enhancing Inference (Sec. 4.5) ], [ Case Study: LLM for OD Calibration (Sec. 4.6) ] ], [ Future Directions and
Challenges (Sec. 5) [ Pitfalls of GenAI for
Transportation Planning (Sec. 5.1) ], [ Pipelines Integrate Transportation
with GenAI (Sec. 5.2) ], [ Data Scarcity and Domain-Specific
Datasets Construction (Sec. 5.3) ], [ Explainability and Hallucination (Sec. 5.4) ], [ Democratizing Access to
Transportation Knowledge (Sec. 5.5) ], [ Call for Novel Evaluation Criteria (Sec. 5.6) ] ], [ Conclusion (Sec. 6) ] ]
Figure 2: Organization of this survey.
2 Background
@@ 2.1 Transportation Planning
2.2 Generative AI Methodologies
2.2.1 Generative AI Models
@@
2.2.2 Core Transportation Applications
@@
Demand Forecasting applies generative AI to predict travel demand patterns across multimodal systems. Models can fine-tune origin-destination (OD) matrices to align with observed traffic counts [109, 7], estimate the adoption of shared mobility services [110, 111, 112], and simulate long-term demand shifts due to demographic or economic changes [70, 52, 113, 88, 114]. These forecasts provide insights into system bottlenecks, enabling planners to balance travel loads across transportation networks effectively.
@@
Benchmark Datasets form the foundation for evaluating generative AI performance in tasks like OD calibration, traffic simulation, and infrastructure planning. Examples include LargeST [124] for traffic analysis and traffic assignment datasets [125]. These datasets provide high-resolution inputs and ground-truth benchmarks to validate AI-driven insights. However, building comprehensive knowledge bases for these models remains a significant challenge. Particularly for forecasting tasks, where uncertainty and dynamic variables play key roles, such as: population growth, land-use changes, and evolving travel behaviors. Generative AI requires diverse, harmonized data inputs, such as historical traffic data, multimodal interactions, and prior model outcomes, but integrating these into a machine-readable format is time-intensive and complex. The lack of standardization, interoperability, and real-time updates further complicates the process, creating a bottleneck in AI adoption. Future advancements must focus on automated dataset synthesis, standardized modeling frameworks, and adaptive real-time updates to ensure that generative AI can address the intricate and evolving demands of transportation systems effectively.
Data Preprocessing Strategies ensure clean, balanced, and representative inputs for AI models. Transportation datasets often require manual or automated annotation of traffic patterns [126], infrastructure attributes [127], or user feedback [128]. Generative AI methods, such as LLM-based automated labeling [129, 130], address annotation challenges while preserving data quality. Data augmentation techniques, such as synthetic OD matrix generation [131, 132], further expand datasets to simulate rare or extreme transportation scenarios.
Fine-Tuning Techniques enable generative AI to adapt to transportation-specific tasks efficiently. Methods like Low-Rank Adaptation (LoRA) [133] optimize model performance while minimizing computational demands. For complex tasks like OD calibration, Chain-of-Thought prompting [115] enhances the model’s reasoning ability to align OD matrices with observed traffic counts.
@@
2.3 Evolution of AI in Transportation
2.3.1 Traditional Approaches
2.3.2 AI-assisted Approaches
2.3.3 GenAI Approaches
The latest evolution, generative AI—represents a paradigm shift toward highly autonomous systems. Generative models leverage large-scale, fine-grained datasets, including real-time sensor inputs and external sources like social media or weather forecasts [69, 70]. These models autonomously generate solutions, simulate traffic scenarios, and optimize infrastructure designs without requiring explicit programming for each task [8, 71]. However, the integration of existing expert knowledge remains crucial in guiding these AI systems. Domain expertise is essential for defining objectives, validating outputs, and ensuring alignment with established transportation modeling principles. Expert insights help refine AI-driven processes by providing a clear direction, especially in complex tasks like multimodal demand forecasting or infrastructure planning, where understanding system nuances is critical. Generative AI systems, when combined with expert knowledge, enable rapid execution and real-time decision-making, accelerating tasks such as traffic rerouting during emergencies, infrastructure design under various constraints, and automated reporting for public engagement.
@@ 3.3 Generative Tasks for Data Synthesis and Scenario Generation
3.4 Simulation Tasks for Traffic Dynamics
@@
  • Computational cost: Another critical challenge is the computational cost associated with running high-resolution, real-time simulations at scale. Generative AI models integrated into digital twins or predictive traffic systems require significant computational resources to process diverse data streams, such as real-time sensor feeds, GPS traces, and dynamic OD matrices [199]. While digital twins offer an opportunity to mirror real-world transportation systems for testing interventions, their reliance on generative AI magnifies the resource demands, particularly for urban-scale or multimodal networks. Developing lightweight and computationally efficient models is crucial to for scalability constraints.
@@ 4.1.1 Benchmark Datasets
4.1.2 Dataset Preparation Strategies
@@
4.1.3 Broad Source of Dataset Collection
@@
Ensuring Data Diversity and Representativeness. To ensure generalizability, transportation datasets must represent diverse geographies, demographic groups, and environmental conditions. For instance, combining urban, suburban, and rural datasets enables generative AI models to address transportation challenges across different spatial contexts. In addition, datasets incorporating varied socioeconomic conditions, such as underserved communities or emerging economies, help mitigate biases and promote equitable solutions.
Data Augmentation for Generative Tasks. Data augmentation techniques enhance dataset size and diversity, especially when real-world data is scarce or incomplete. For traffic forecasting, techniques such as synthetic trajectory generation and perturbation modeling can create realistic yet varied travel patterns. In infrastructure optimization, generative models can simulate alternative urban layouts, enabling planners to evaluate multiple design scenarios under different constraints, such as cost, population density, and environmental impact.
@@
4.2 Fine-Tuning Generative AI for Transportation Planning
This section presents the general technique of fine-tuning generative AI models, then demonstrates how the technique is adopted in the specific context of OD Demand Calibration in transportation planning.
4.2.1 General Steps in Fine-Tuning Generative AI Models
Fine-tuning refers to adapting a pre-trained generative model to better suit a specific downstream task. The process generally involves:
@@ 4.2.1 General Steps in Fine-Tuning Generative AI Models
    Data Preprocessing: Aligning input-output data pairs through normalization, formatting, and noise filtering to create structured training examples.
@@
    Model Adaptation: Using parameter-efficient methods such as LoRA or prefix-tuning to modify a small subset of weights.
@@
    Training: Performing loss minimization on structured data using strategies like gradient accumulation and mixed-precision training, while validating on held-out subsets.
@@
    Prompt Engineering: Designing task-specific prompts to condition the model outputs in alignment with domain requirements.
  5. 5.
    Evaluation: Measuring performance using relevant metrics such as RE (Relative Error) and RMSE.
4.2.2 Specific Problem Setting: OD Demand Calibration
OD (Origin-Destination) Demand Calibration estimates travel demand between regions by adjusting OD matrices to match observed traffic counts. Accurate OD matrices are essential for effective infrastructure planning and traffic management.
Domain-Specific Dataset
The OpenTI dataset [7] supports OD calibration with rich traffic flows, network structures, and travel demands. It provides:
  • Input: Raw OD matrices and observed link-level traffic counts.
  • Output: Calibrated OD matrices aligned with observed traffic patterns across time intervals.
Adopting Fine-Tuning Steps for OD Calibration
Each general step is instantiated in this application as follows:
1. Data Preprocessing: Preprocessing OpenTI data ensures consistency between inputs (raw OD matrices and observed traffic counts) and outputs (calibrated OD matrices). The steps include: (1) Normalization: Scaling OD demand values to ensure uniform ranges across datasets. (2) Matrix Formatting: Aligning OD matrices with observed traffic counts to maintain input-output coherence. (3) Noise Reduction: Filtering out inconsistencies or redundant data points to streamline the training process. The preprocessed data pairs provide well-structured examples for training generative AI models to produce calibrated OD matrices.
2. Model Adaptation: Apply LoRA [133] or prefix-tuning [267] to an LLM such as Llama3.1-8B.
3. Training Process: Utilize GPU clusters, gradient accumulation [268], and mixed-precision training [269]. Minimize loss between generated and ground-truth calibrated OD matrices.
4. Prompt Engineering: Craft prompts that direct the model to adjust demand values while preserving traffic consistency:
@@ 4.2.2 Specific Problem Setting: OD Demand Calibration
5. Evaluation:
  • Relative Error (RE): Compares predicted vs observed counts.
  • RMSE: Measures deviation from ground-truth OD matrices.
Expected Outputs and Impact
After fine-tuning on the OpenTI dataset, the generative AI model is expected to perform effectively across several OD calibration tasks:
@@ 5. Evaluation:
4.3 Zero-Shot Inference with Generative AI for Transportation Accessibility
This section presents the general framework of Zero-Shot Learning (ZSL) using generative AI, and demonstrates how it is applied to estimating city-level public transportation accessibility, particularly through People Near Transit (PNT) scores.
4.3.1 General Methodology: Zero-Shot Learning (ZSL) with LLMs
ZSL enables a pre-trained LLM to perform domain-specific tasks without any fine-tuning or labeled examples. This is achieved solely through carefully constructed prompts. For urban and transportation studies, ZSL offers a scalable and cost-effective tool for analyzing cities lacking structured datasets. Key Features of ZSL are: (1) No training required: Tasks are performed via prompt design without additional data or tuning. (2) Wide applicability: General knowledge encoded during pretraining is leveraged for domain-specific inference. (3) Interpretability: Prompts and outputs are transparent and human-readable, supporting validation.
4.3.2 Specific Problem Setting: Estimating PNT in Global Cities
People Near Transit (PNT) measures the proportion of a city’s population living within a defined distance (e.g., 500m, 1000m, 1500m) of transit stops. Directly obtaining such data requires detailed GIS and fieldwork, which are often unavailable.
Input Representation
To approximate PNT using LLMs, we convert spatial accessibility into qualitative prompts about urban conditions, including urban planning quality, road network conditions, public transportation extensiveness, and government support for transit development. Each factor is queried using a 0.0 to 10.0 scale, allowing the LLM to infer high-level characteristics of accessibility.
Adapting ZSL for Transportation Accessibility
The most important step in applying ZSL is prompt engineering, i.e., Design prompts incorporating contextual information. An example prompt used in [275] is shown below.
Outputs and Impact
Compare inferred responses with ground-truth PNT scores, where available, across spatial scales (500m, 1000m, 1500m). The responses generated by LLMs can be validated by directly correlating their outputs with ground-truth PNT (Proximity to Network Transit) scores [272], where available. This approach provides a robust means of assessing the accuracy of predictions. Notably, this method demonstrates significant improvements compared to directly querying the LLMs for PNT scores without further calibration. Table 2 summarizes the performance of LLM-generated predictions across varying spatial scales (500m, 1000m, and 1500m) in 85 global cities. The comparison covers three methods: Zero-Shot Learning (ZSL), direct queries, and using average PNT scores as predictions (AVG), and we find that ZSL consistently performs better than traditional methods. For cities without available transit data, the LLM’s estimations provide a scalable and efficient approximation, which can support decision-making and hypothesis testing.